Today 7:03 AM Details Question Completion Status: | « | | Question 2 of 29 > Moving to another question will save this response Question 2 0.5 points Save Answer A food cart sells Indian food (Corys favorite!). On average, 25 customers arrive per hour (randomly distributed throughout the hour, much like real life). It is able to serve 30 customers per hour. The food cart pays one server $15 an hour and estimates that the cost of a customer waiting in line for an hour is $12 (from lost revenue of reneging, reputational damage, etc.). What is the cost of servers plus the cost of waiting per hour? Question 2 of 29
media%2Fc81%2Fc8166ff6-8759-4b7d-a5b9-8a
media%2F7ab%2F7abc20fb-8a3b-44c6-805f-cf
media%2Fe1c%2Fe1cfea3b-0ef5-4f91-a0ad-2c
media%2Feb7%2Feb7360ef-a1c6-498e-8785-b1
media%2Fc38%2Fc3827fea-e5df-4ad2-be1e-d1
Answer #1

Q2)

Arrival rate, a = 25 customers per hour

Service rate, s = 30 customers per hour

Server cost, Cs = $ 15 per hour

Waiting cost, Cw = $ 12 per hour

Average number of customers waiting, Lq = a2/(s*(s-a) = 252/(30*(30-25) = 4.1667

Cost of server plus cost of waiting per hour = Cs + Lq*Cw = 15 + 4.1667*12 = $ 65

Q3) False

Producing in-house requires a fixed cost, where purchasing from external supplier does not require a fixed cost. Therefore, for lower volumes, it is cheaper to buy from external supplier than make in-house

Q6) EOQ = SQRT(2DS/H) = SQRT(2*1000*375/300) = 50

Total inventory costs per year = (D/Q)*S + (Q/2)*H = (1000/50)*375 + (50/2)*300 = $ 15,000

Q9) Average time a customer spend in entire system, W = 1/(s-a) = 1/(30-25) = 0.2 hour or 12 minutes

Q13) Optimal order quantity (EOQ) = SQRT(2*20000*25/0.25) = 2000

Q23) Number of customers in system, L = W*a = 0.2*25 = 5 customers